OpenStudy (anonymous):

I am trying to determine the rate which a human body drops to the bottom of the ocean. I've found two different versions of this equation, but I'm unsure of the results they yield, or if I should be using something else. The first one is: R = (Sg (D1-D2))/µ where R=sinking rate; S=cross-sectional area of object; g=gravitational acceleration; D1=density of the object; D2=density of fluid in the environment; µ=dynamic viscosity The second one is: U = [2(a^2)g(p-p0)] / 9µ where U=sinking rate; a=radius, g=gravitational constant; p=density of organism, po=density of medium; m=dynamic viscosity S=1.61m^2 based on height (5'2") & weight (130lbs) of the person density of human body=1010kg/m^3 density of saltwater=1025kg/m^3 dynamic viscosity=0.001674kg/ms Saltwater temp=29F/2C

OpenStudy (anonymous):

Are you planning to drop a body to bottom of the ocean?.....I don't think I want to help with this no matter what your answer is. You gonna have to go alone on this one in case u a cearal killer.

OpenStudy (mos1635):

2nd formula seams to aplay ty sfairical objects (a=radious) so i would try the 1st one

OpenStudy (anonymous):

This is to investigate a scuba diving death where a witness said the person dropped at a very fast rate all the way to 187ft from a depth of about 20ft. The diver weighted 130lbs and had about 60lbs of additional weight. I want to find out if this is possible or if something else played a part in the sudden descent. I solved the first equation but it gave me a drop rate of 141524 m/s which doesn't make sense whatsoever. Thus my need for help in this matter.

OpenStudy (mos1635):

 A=D2*V*g W=m*g=D1*V*g F=μ*S*v (or something like that) terminal velocity when ΣF=0 W-A-F=0 F=W-A v=(D1-D2)*V*g/μ*S that is the formula i derived. look's like the one you stated but it"not. assumind volume of dicer V=0.3*0.3*1.8=0.163m^3 i have v=11.7 m/sec note: dynamic viscosity=0.001674kg/ms=1.674 Kg/s

OpenStudy (anonymous):

I see what I did wrong, I thought that the ms in the viscosity constant was meters per seconds, not miliseconds. I found an equation for calculating the volume for an adult human being that claims to be 3.5% within the true value. V = S{[51.44(W/H)]+15.3}, where V is in L, S in m^2, W in kg and H in cm.